# A Unified Representation of Neural Networks Architectures

### v1

@@ Abstract
In this paper we consider the limiting case of neural networks (NNs) architectures when the number of neurons in each hidden layer and the number of hidden layers tend to infinity thus forming a continuum, and we derive approximation errors as a function of the number of neurons and/or hidden layers. Firstly, we consider the case of neural networks with a single hidden layer and we derive an integral infinite width neural representation that generalizes existing continuous neural networks (CNNs) representations. Then we extend this to deep residual CNNs that have a finite number of integral hidden layers and residual connections. Secondly, we revisit the relation between neural ODEs and deep residual NNs and we formalize approximation errors via discretization techniques. Then, we merge these two approaches into a unified homogeneous representation of NNs as a Distributed Parameter neural Network (DiPaNet) and we show that most of the existing finite and infinite-dimensional NNs architectures are related via homogeneization/discretization with the DiPaNet representation. Our approach is purely deterministic and applies to general, uniformly continuous matrix weight functions. Differences and similarities with neural fields are discussed along with further possible generalizations and applications of the DiPaNet framework.
@@
Neural networks, Residual neural networks, Neural ODEs, Distributed parameter systems
@@ 1 Introduction
Another approach to designing continuous neural networks stems from the neuro-physiological perspective that treats neurons as entities distributed in space, see, e.g., (Amari2000Methods; Cook2021NeuralFieldOverview). Neural fields can be regarded as neural PDEs where one integrates over the space of neurons. Such neural field representations have been recently employed to develop the Neural Integro-Differential Equations (NIDEs) framework (Zappala2023NeuralIntegroDifferential) for modeling dynamics with integral-differential equations that are determined by non-local operators. Furthermore, delayed neural fields have been used for observer and controller design for PDEs in (Brivadis2024AdaptiveObserver). Although neural operators (Kovachki2023) and NIDEs (Zappala2023NeuralIntegroDifferential) represent generalizations of neural ODEs (neuralODE) to functions spaces, a neural network architecture that unifies integral continuous neural networks for function approximation as introduced by (cnn) with neural ODEs is yet to be developed, to the best knowledge of the authors.
In this work we thus develop a novel homogeneous representation of NNs that unifies the CNN representation of (cnn) with the neural ODE representation of (neuralODE) via a distributed parameter systems approach. Our framework is fully deterministic and yields a neural integral differential representation where the integration is over an arbitrary domain, e.g., $[0,1)$. The resulting homogeneous neural network can be regarded as a Distributed Parameter neural Network (DiPaNet) from which we can recover neural ODEs (neuralODE), Deep CNNs (cnn), and Deep Residual NNs (he2016deep; Tabuada) by applying suitable discretization techniques. In reverse, we can show that all these architectures converge to the DiPaNet representation through suitable homogenization techniques.
@@ 2.1 DeepNet and DeepCNN
The last part of this definition deals with the increments of consecutive terms. Intuitively, a sequence of hyperparameters is of bounded variation if the sum of all increments is bounded. This is needed to define the homogeneization part of our first main result.
@@
$\bullet$ Conversely the homogeneization of a DeepNet as given by (1), as $n$ tends to infinity, is a DeepCNN modeled by (3). More specifically, given consistent sequences of matrices $(L^{n})_{n>0}$, $(P^{n})_{n>0}$, $(W^{n}_{i})_{n>0}$ with bounded variation, a consistent set of activation functions $(\sigma^{n}_{i})_{n>0}$ with bounded variation, for $i=1,\ldots,\ell$, and positive values $r$ and $\varepsilon$, there exist functions $\mathbf{L}$, $\mathbf{P}$, $\mathbf{W}_{i}$, and $\boldsymbol{\sigma}_{i}$, for $i=1,\ldots,{\ell}$, and a sufficiently large integer $\overline{n}$ satisfying the property, for all $X\in\mathbb{R}^{p}$ satisfying $\|X\|\leq r$, the distance between the outputs of (1) with $n\geq\overline{n}$ neurons and of (3) is less than $\varepsilon$:
@@ 2.2 DeepResNet and DeepResCNN
$\bullet$ The homogeneization of any sequence of DeepResNet as given by (9) (with consistent sequences of matrices) is a DeepCNN when the number of neurons increase and go to infinity.
@@ 3.1 DeepResNet and NeuralODE
$\bullet$ Conversely the homogeneization of a DeepResNet as given by (8), as $\ell$ tends to infinity, is a NeuralODE modeled by (12). More specifically, given matrices $L\in\mathbb{R}^{n\times p}$, $P\in\mathbb{R}^{q\times n}$, an uniformly bounded and equicontinuous sequence of matrices $((W^{\ell}_{i})_{i=1,\ldots,\ell})_{\ell>0}$, an uniformly bounded and equicontinuous sequence of functions $((\ell\sigma^{\ell}_{i})_{i=1,\ldots,\ell})_{\ell>0}$ and positive values $r$ and $\varepsilon$.
@@ 4.1 DeepResNet and DiPaNet
A neural integro-differential equation similar to (15) was also derived in (Zappala2023NeuralIntegroDifferential), albeit with a different purpose, i.e., modeling of non-local dynamics represented by integro-differential equations, and with a different construction of an integral kernel operator. Furthermore, (Zappala2023NeuralIntegroDifferential) did not establish explicit connections with classical NNs such as DeepResNets and other NNs as considered in Sections 2-3 of this paper.
The above-defined class of NNs is a distributed parameter system because the state $Z$ depends on the time, and of functions depending themselves on both variables $t$ and $\tau$. We call such a NN as a Distributed Parameter neural Network (DiPaNet). The links of this NN with respect to the other ones previously studied in this paper lie in the discretization (or conversely in homogeneization) in width. This is performed via standard numerical discretization (or conversely by interpolation) in the domain $[0,1)$ to approximate the integral from 0 to $1$ in (15), as done in the proof of Theorem 3 when dealing with DeepCNN. Another link with one previously considered NN is the discretization (or conversely homogeneization) in depth that is performed via the dynamical system perspective in (15), as done in the proof of Theorem 7 dealing with NeuralODE, rather than numerical discretization of the integral from 0 to $T$ in (14) for example.
@@
$\bullet$ The homogeneization of a DeepResCNN as given by (10), as $\ell$ tends to infinity, is a DiPaNet modeled by (15).
@@
$\bullet$ The homogeneization of a NeuralODE as given by (11), as $n$ tends to infinity, is a DiPaNet modeled by (15).
@@
$\bullet$ The homogeneization of a DeepResNet as given by (8), as $\ell$ and $n$ tend to infinity, is a DiPaNet modeled by (15).

### v2

@@ Abstract
In this paper we consider the limiting case of neural networks (NNs) architectures when the number of neurons in each hidden layer and the number of hidden layers tend to infinity thus forming a continuum, and we derive approximation errors as a function of the number of neurons and/or hidden layers. Firstly, we consider the case of neural networks with a single hidden layer and we derive an integral infinite width neural representation that generalizes existing continuous neural networks (CNNs) representations. Then we extend this to deep residual CNNs that have a finite number of integral hidden layers and residual connections. Secondly, we revisit the relation between neural ODEs and deep residual NNs and we formalize approximation errors via discretization techniques. Then, we merge these two approaches into a unified homogeneous representation of NNs as a Distributed Parameter neural Network (DiPaNet) and we show that most of the existing finite and infinite-dimensional NNs architectures are related via homogenization/discretization with the DiPaNet representation. Our approach is purely deterministic and applies to general, uniformly continuous matrix weight functions. Relations with neural fields and other neural integro-differential equations are discussed along with further possible generalizations and applications of the DiPaNet framework.
@@
Neural networks, Continuous neural networks, Residual neural networks, Neural ODEs, Distributed parameter systems
@@ 1 Introduction
Another approach to designing continuous neural networks, i.e., the neural fields approach, stems from the neuro-physiological perspective that treats neurons as entities distributed in space, see, e.g., (Amari2000Methods; CoombesEtAl2014; Cook2021NeuralFieldOverview). Neural fields can be regarded as neural PDEs where one integrates over the space of neurons. Such neural field representations have been recently employed to develop the Neural Integro-Differential Equations (NIDEs) framework (Zappala2023NeuralIntegroDifferential) for modeling dynamics with integro-differential equations that are determined by non-local operators. Furthermore, delayed neural fields have been used for observer and controller design for PDEs in (Brivadis2024AdaptiveObserver). It is worth to mention that existing neural fields and neural integro-differential representations have been motivated by learning dynamics described by partial differential equations (PDEs) and especially brain modeling problems.
Although neural operators (Kovachki2023) and NIDEs (Zappala2023NeuralIntegroDifferential) represent generalizations of neural ODEs (neuralODE) to functions spaces, a neural network architecture that unifies integral continuous neural networks for function approximation as introduced by (cnn) with neural ODEs is yet to be developed, to the best knowledge of the authors.
In this paper we analyze the limit of deep residual neural networks when the number of neurons and the number of hidden layers tend to infinity and we derive a novel homogeneous representation of NNs that unifies the CNN representation of (cnn) with the neural ODE representation of (neuralODE) via a distributed parameter systems approach. Our framework is fully deterministic and yields a neural integro-differential representation where the integration is over an arbitrary domain, e.g., $[0,1)$. The resulting homogeneous neural network can be regarded as a Distributed Parameter neural Network (DiPaNet) from which we can recover neural ODEs (neuralODE), Deep CNNs (cnn), and Deep Residual NNs (he2016deep; Tabuada) by applying suitable discretization techniques. In reverse, we can show that all these architectures converge to the DiPaNet representation through suitable homogenization techniques.
@@ 2.1 DeepNet and DeepCNN
The last part of this definition deals with the increments of consecutive terms. Intuitively, a sequence of hyperparameters is of bounded variation if the sum of all increments is bounded. This is needed to define the homogenization part of our first main result.
@@
$\bullet$ Conversely the homogenization of a DeepNet as given by (1), as $n$ tends to infinity, is a DeepCNN modeled by (3). More specifically, given consistent sequences of matrices $(L^{n})_{n>0}$, $(P^{n})_{n>0}$, $(W^{n}_{i})_{n>0}$ with bounded variation, a consistent set of activation functions $(\sigma^{n}_{i})_{n>0}$ with bounded variation, for $i=1,\ldots,\ell$, and positive values $r$ and $\varepsilon$, there exist functions $\mathbf{L}$, $\mathbf{P}$, $\mathbf{W}_{i}$, and $\boldsymbol{\sigma}_{i}$, for $i=1,\ldots,{\ell}$, and a sufficiently large integer $\overline{n}$ satisfying the property, for all $X\in\mathbb{R}^{p}$ satisfying $\|X\|\leq r$, the distance between the outputs of (1) with $n\geq\overline{n}$ neurons and of (3) is less than $\varepsilon$:
@@ 2.2 DeepResNet and DeepResCNN
$\bullet$ The homogenization of any sequence of DeepResNet as given by (9) (with consistent sequences of matrices) is a DeepCNN when the number of neurons increase and go to infinity.
@@ 3.1 DeepResNet and NeuralODE
$\bullet$ Conversely the homogenization of a DeepResNet as given by (8), as $\ell$ tends to infinity, is a NeuralODE modeled by (12). More specifically, given matrices $L\in\mathbb{R}^{n\times p}$, $P\in\mathbb{R}^{q\times n}$, an uniformly bounded and equicontinuous sequence of matrices $((W^{\ell}_{i})_{i=1,\ldots,\ell})_{\ell>0}$, an uniformly bounded and equicontinuous sequence of functions $((\ell\sigma^{\ell}_{i})_{i=1,\ldots,\ell})_{\ell>0}$ and positive values $r$ and $\varepsilon$.
@@ 4.1 DeepResNet and DiPaNet
A neural integro-differential equation similar to (15) was also derived in (Zappala2023NeuralIntegroDifferential), albeit with a different purpose, i.e., modeling of non-local dynamics represented by integro-differential equations, and with a different construction of an integral kernel operator. Furthermore, (Zappala2023NeuralIntegroDifferential) did not establish formal relations with classical NNs such as DeepResNets and other NNs as considered in Sections 2-3 of this paper.
The above-defined class of NNs is a distributed parameter system because the state $Z$ depends on the time, and of functions depending themselves on both variables $t$ and $\tau$. We call such a NN as a Distributed Parameter neural Network (DiPaNet). The links of this NN with respect to the other ones previously studied in this paper lie in the discretization (or conversely in homogenization) in width. This is performed via standard numerical discretization (or conversely by interpolation) in the domain $[0,1)$ to approximate the integral from 0 to $1$ in (15), as done in the proof of Theorem 3 when dealing with DeepCNN. Another link with one previously considered NN is the discretization (or conversely homogenization) in depth that is performed via the dynamical system perspective in (15), as done in the proof of Theorem 7 dealing with NeuralODE, rather than numerical discretization of the integral from 0 to $T$ in (14) for example.
@@
$\bullet$ The homogenization of a DeepResCNN as given by (10), as $\ell$ tends to infinity, is a DiPaNet modeled by (15).
@@
$\bullet$ The homogenization of a NeuralODE as given by (11), as $n$ tends to infinity, is a DiPaNet modeled by (15).
@@
$\bullet$ The homogenization of a DeepResNet as given by (8), as $\ell$ and $n$ tend to infinity, is a DiPaNet modeled by (15).